4/5+0.25(decimal answer) the possibilites are..........0.65, 1.05, 21/20, or none of these
Answer by DWL(56)

Convert 4/5 to decimal form:
4/5 = .80
.80 + .25 = 1.05
Answer = 1.05
or
Convert .25 to fraction; .25 = 1/4
then:
4/5 + 1/4
Find common denominator, in this case 20 is a number divisable by both 5 and 4.
4/5 converts to 16/20 ((20/5)*4) and 1/4 to 5/20 ((20/4)*1)
Now add 16/20 + 5/20 = 21/20
Answer can also be 21/20